OpenStudy (anonymous):

WILL MEDAL AND FAN! PLEASE, THIS IS URGENT! Solve √(k-7) = √(k) -1 for k.

Nnesha (nnesha):

take square both sides to cancel out square root

OpenStudy (anonymous):

i tried that and i got this \[k-7= k-1\]

OpenStudy (anonymous):

or is it k+1? i think its K-1

OpenStudy (anonymous):

can anyone help me please?

OpenStudy (freckles):

\[\sqrt{k-7}=\sqrt{k}-1 ?\]

OpenStudy (anonymous):

I'm trying to have it done by the time my best friend gets on.

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

the choices are: 20 16 4 8

Nnesha (nnesha):

oh nvm I didn't notice the parentheses

OpenStudy (freckles):

square both sides

OpenStudy (freckles):

you will need this on the right hand side \[(a-b)^2=a^2+b^2-2ab\]

OpenStudy (anonymous):

\[\sqrt{k-7} = \sqrt{k} -1\]

OpenStudy (freckles):

have you expanded the right hand side yet?

OpenStudy (anonymous):

i dont know what that means

OpenStudy (anonymous):

i got to this point\[k-7 = k-1\]

OpenStudy (freckles):

\[\sqrt{k-7}=\sqrt{k}-1 \\ \text{ Square both sides } (\sqrt{k-7})^2=(\sqrt{k}-1)^2 \\ k-7=(\sqrt{k}-1)^2\] recall \[(a-b)^2=a^2+b^2-2ab\] you can use this to expand the square part on the right hand side

OpenStudy (anonymous):

ooooohhhhhhh soooooo\[k+1-2k\]

OpenStudy (anonymous):

right? so \[k+1-2k=k-7\]

OpenStudy (freckles):

close

OpenStudy (freckles):

you had a sqrt( ) on k

OpenStudy (freckles):

\[(a-b)^2=a^2+b^2-2ab \\ (\sqrt{k}-1)^2=(\sqrt{k})^2 +(1)^2-2(\sqrt{k})(1) \\ (\sqrt{k}-1)^2=k+1-2 \sqrt{k}\]

OpenStudy (anonymous):

oh right. \[k-7 = k+1 - 2\sqrt{k}\]

OpenStudy (freckles):

so you really have \[k-7=k+1-2 \sqrt{k} \text{ or subtracting } k \text{ on both sides } -7=1-2 \sqrt{k}\]

OpenStudy (freckles):

isolate the sqrt(k) part

OpenStudy (anonymous):

k\[-2\sqrt{k}=-8\]

OpenStudy (anonymous):

right?

OpenStudy (freckles):

yeah

OpenStudy (freckles):

divide both sides by -2

OpenStudy (anonymous):

\[\sqrt{k}=4\]

OpenStudy (anonymous):

k=16!

OpenStudy (freckles):

yep and you can check that by pluggin it in

OpenStudy (anonymous):

THANK YOUUU soooooooooo muchhh

OpenStudy (anonymous):

Im a fan a medals

OpenStudy (freckles):

\[\sqrt{k-7}=\sqrt{k}-1 \\ k=16 \\ \sqrt{16-7}=\sqrt{9}=3 \\ \text{ and the other side } \sqrt{16}-1=4-1=3 \] 3=3 k=16 is definitely a solution since it gives a true equation when pluggin it back in

OpenStudy (anonymous):

thank youuu

OpenStudy (freckles):

np